### v1

# Diffusion Models Meet Remote Sensing: Principles, Methods, and Perspectives

Abstract
As a newly emerging advance in deep generative models, diffusion models have achieved state-of-the-art results in many fields, including computer vision, natural language processing, and molecule design. The remote sensing community has also noticed the powerful ability of diffusion models and quickly applied them to a variety of tasks for image processing. Given the rapid increase in research on diffusion models in the field of remote sensing, it is necessary to conduct a comprehensive review of existing diffusion model-based remote sensing papers, to help researchers recognize the potential of diffusion models and provide some directions for further exploration. Specifically, this paper first introduces the theoretical background of diffusion models, and then systematically reviews the applications of diffusion models in remote sensing, including image generation, enhancement, and interpretation. Finally, the limitations of existing remote sensing diffusion models and worthy research directions for further exploration are discussed and summarized.
Index Terms:
Diffusion Models, Remote Sensing, Generative Models, Deep Learning.
I Introduction
Remote sensing (RS), as an advanced earth observation technology, has been widely used in civilian and military fields such as environmental monitoring, urban planning, disaster response, and camouflage detection [1, 2, 3, 4, 5]. Following the boom of artificial intelligence, employing deep learning models to interpret RS images has become a large-scale solution for these applications [6]. Early intelligent RS interpretation methods primarily relied on supervised deep neural networks, which were trained with massive data and high-quality annotations. However, the scarcity of annotations and the high acquisition costs of RS images have hindered further advancements in these methods.
Figure 1: Development of diffusion models in RS. Statistical data as of the first quarter of 2024.
The advent of deep generative models effectively solves the problems in the above supervised interpretation methods, bringing new opportunities for the intelligent processing of RS images. Specifically, deep generative models are capable of learning the data distribution from limited RS images to generate new data samples. At the same time, they can generate annotation information directly from low-quality or unlabeled RS images by learning the mapping relationships between images, reducing the need for high-quality manual annotations. Furthermore, they have also demonstrated excellent capabilities in learning representations for image details and complex scenes. Over the past decade, numerous works on deep generative models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Normalizing Flows (NFs), have sprung up to tackle the challenges associated with RS images [7, 8, 9, 10, 11]. Despite the flourishing development of these generative models, each comes with its limitations. For example, VAE [12] requires a trade-off between reconstruction loss (similarity between the output and the input) and latent loss (proximity of the hidden nodes to the normal distribution), so that the generated images are often blurry. The structure of NF-based models [13] needs to comply with the calculation of probability density, resulting in limited scalability and flexibility. While GAN [14] has many variants and generates high-quality RS images, its training process is unstable and easy to collapse (i.e., the generated sample pattern is singular and cannot cover diverse patterns).
Figure 2: The training procedure of denoising diffusion probabilistic model (DDPM), where yellow lines represent the forward diffusion process, and blue lines represent the backward diffusion process.
In this context, diffusion models [15], as a newly emerging type of deep generative models, have brought about a revolutionary advancement in artificial intelligence. By modeling the inverse process of transforming regular images into random noise, diffusion models have demonstrated unprecedented performance in 2D and 3D image generation [16, 17, 18, 19, 20], image editing [21, 22, 23, 24], image translation [25, 26, 27, 28], and other computer vision tasks [29, 30, 31, 32]. Moreover, they have achieved state-of-the-art results in many other fields, including natural language processing [33, 34, 35], audio synthesis [36, 37, 38], and molecular design [39, 40, 41], challenging the long-standing dominance of GANs.
Given these remarkable achievements, the RS community has also quickly applied diffusion models to a variety of tasks for image processing. Since 2021, the application of diffusion models in RS has shown a rapid development trend of expanding scope and increasing quantity (see Fig. 1). In fact, diffusion models have significant advantages over other deep generative models in processing and analyzing RS images.
  • Firstly, due to the atmospheric interference and limitations of imaging equipment, RS images often contain noise. The inherent denoising ability of diffusion models can rightly eliminate these negative effects.
  • Secondly, RS images are highly diverse due to differences in collection time, equipment, and environment. The architecture of diffusion models is flexible, allowing the introduction of conditional constraints to cope with various changes.
  • Thirdly, RS images always contain diverse and complex scenes. The precise mathematical derivation and progressive learning process of diffusion models offer advantageous in learning such complex data distribution.
  • Additionally, diffusion models can provide more stable training than GANs, which is suitable for training large-scale RS datasets.
In summary, diffusion models possess great development potential in the field of RS. Therefore, it is necessary to review and summarize existing diffusion model-based RS papers to help researchers gain a comprehensive understanding of current research status, and identify the gaps in the application of diffusion models in RS, thereby promoting further development in this field.
The remainder of this paper is organized as follows. Section II introduces the theoretical background of diffusion models. Section III reviews the application of diffusion models across various RS image processing tasks, and demonstrates the superiority of diffusion models through a series of visual experimental results and quantitative metrics. Section IV discusses the limitations of the existing RS diffusion models and reveals possible research directions in the future. Finally, conclusions are drawn in Section V.
II Theoretical Background of Diffusion Models
Diffusion models, also known as diffusion probabilistic models, are a family of deep generative models. In general, generative model is to convert a random distribution (i.e., noise) into a “probability distributions” that is the same as the distribution of the observed dataset, and obtain the desired outcomes by sampling from this “probability distribution”.
Obviously, it is quite difficult to obtain the target distribution directly, but disrupting a regular distribution into random noise is straightforward, and can be achieved by continuously adding Gaussian noise, as illustrated in Fig. 2. Diffusion model is inspired by this process and realizes the generation of target data by learning the reverse denoising process of this process, which can be traced back to 2015 [42] and became popular after denoising diffusion probabilistic model (DDPM) was published in 2020 [15].
II-A Denoising Diffusion Probabilistic Model (DDPM)
As shown in Fig. 2, the training procedure involves two phases: the forward diffusion process and the backward diffusion process.
Forward Diffusion Process: Given the original image $x_{0}$, this process generates the noise-contaminated images $x_{1},x_{2},\ldots,x_{T}$ through $T$ iterations of noise addition. where the image $x_{t}$ obtained at each step is only related to $x_{t-1}$. Thus, this process can be represented by a Markov chain:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (1)
$q(x_{1:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1})=\prod_{t=1}^{T}\mathcal{N}(x_{t};\sqrt{1-\beta_{t-1}},\beta_{t}I)$ (2)
where $q(x_{t}|x_{t-1})$ is the transition probability of the Markov chain, which represents the distribution of Gaussian noise added at each step. $\beta_{t}$ is a hyperparameter for the variance of the Gaussian distribution, linearly increasing with $t$. $I$ denotes the identity matrix with the same dimensions as the input image $x_{0}$.
An important property of the forward process is that it allows to directly obtain any noised image $x_{t}$ from the original image $x_{0}$ and $\beta_{t}$, which is achieved through the reparameterization technique [12]. Specifically, with the notation of $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, Eq. (1) could be expanded as
$x_{t}$ $=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1}$
$=\sqrt{\alpha_{t}}\left(\sqrt{\alpha_{t-1}}x_{t-2}+\sqrt{1-\alpha_{t-1}}\epsilon_{t-2}\right)+\sqrt{1-\alpha}_{t}\epsilon_{t-1}$
$=\sqrt{\alpha_{t}\alpha_{t-1}}x_{t-2}+\sqrt{1-\alpha_{t-1}\alpha_{t-2}}\bar{\epsilon}_{t-2}$
=$\ldots$= …
$=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\bar{\epsilon}$ (3)
where $\epsilon_{t}$, $\epsilon_{t-1}\sim\mathcal{N}(0,I)$ and $\bar{\epsilon}_{t-2}$ is their merged result. According to the additivity of the independent Gaussian distribution, i.e., $\mathcal{N}(0,\sigma_{1}^{2}I)+\mathcal{N}(0,\sigma_{2}^{2}I)\sim\mathcal{N}(0,(\sigma_{1}^{2}+\sigma_{2}^{2})I)$, the third line of Eq. (3) conforms to Gaussian distribution, which means that the final derivation result also conforms to Gaussian distribution. Therefore, any noised image $x_{t}$ satisfies:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},1-\bar{\alpha}_{t}I)$ (4)
In this way, when $T\to\infty$, $x_{T}$ can converge to the standard normal distribution $\mathcal{N}(0,I)$, consistent with the original design intention.
Backward Diffusion Process: This process aims to obtain the reversed transition probability $q(x_{t-1}|x_{t})$, thereby gradually restoring the image $\hat{x_{0}}$ from the noise. However, $q(x_{t-1}|x_{t})$ is difficult to solve explicitly, so a neural network is employed to learn this distribution.
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (5)
where $\theta$ represents the parameters of the neural network to be optimized, and the network is typically based on a U-Net architecture [43]. Accordingly, the backward diffusion process can be expressed as
$p_{\theta}(x_{0:T})$ $=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t})$ (6)
The training goal of the network is to match the backward diffusion process $p_{\theta}(x_{0},x_{1},\ldots,x_{T})$ with the forward diffusion process $p_{\theta}(x_{0},x_{1},\ldots,x_{T})$ , which can be achieved by minimizing the Kullback-Leibler (KL) divergence:
${{\cal L}}(\theta)$ $=\text{KL}(q(x_{0},x_{1},\ldots,x_{T})||p_{\theta}(x_{0},x_{1},\ldots,x_{T}))$
$=-\mathbb{E}_{q(x_{0:T})}[\log p_{\theta}(x_{0},x_{1},\ldots,x_{T})]+\text{const}$ (7)
$=-\mathbb{E}_{q(x_{0:T})}[-\log p(x_{T})-\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t}|x_{t-1})}]+\text{const}$
Here, $const$ denotes a constant independent of $\theta$, and the first term of Eq. (7) represents the variational lower-bound of the negative log-likelihood, similar to VAE.
Notably, when the prior $x_{0}$ is introduced in $q(x_{t-1}|x_{t})$, it can be converted by Bayes Rule
$q(x_{t-1}|x_{t},x_{0})$ $=\frac{q(x_{t},x_{0},x_{t-1})}{q(x_{t},x_{0})}$
$=\frac{q(x_{0})q(x_{t-1}|x_{0})q(x_{t}|x_{t-1},x_{0})}{q(x_{0})q(x_{t}|x_{0})}$
$=q(x_{t}|x_{t-1},x_{0})\frac{q(x_{t-1}|x_{0})}{q(x_{t}|x_{0})}$ (8)
where $q(x_{t}|x_{t-1},x_{0})$ is defined in Eq. (1), $q(x_{t-1}|x_{0})$ and $q(x_{t}|x_{0})$ can be obtained by Eq. (4). After simplification, Eq. (8) can be rewritten as
$q(x_{t-1}|x_{t},x_{0})=\mathcal{N}(x_{t-1};\tilde{\mu}_{t}(x_{t}),\tilde{\beta}_{t}\mathbf{I})$ (9)
$\tilde{\mu}_{t}(x_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\alpha_{t}}}\bar{\epsilon}\right)$ (10)
$\hat{\beta}_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ (11)
$\alpha_{t}$ and $\beta_{t}$ are both constants in the above equation, only $\bar{\epsilon}$ in Eq. (10) can be parameterized by the neural network as
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)$ (12)
In other words, the constructed neural network is to learn the noise $\epsilon_{\theta}(x_{t},t)$, which is reasonable since the process from $x_{t}$ to $x_{t-1}$ is essentially a denoising process.
According to [15], the optimization goal of the network can be further simplified with the help of Eq. (12) and (3) as the following form:
${{\cal L}}_{simple}(\theta)=\mathbb{E}_{x_{0},\epsilon,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right]$ (13)
which intuitively shows that the core of the diffusion model is to minimize the distance between the predicted noise $\epsilon_{\theta}$ and the actual noise $\epsilon$.
Figure 3: The sampling process of DDPM. Supposing that sampling begins at T=1000, the noise distribution $\epsilon_{\theta}(y_{t},t)$ is obtained from the well-trained diffusion model. Then, the noised image $Y_{t}$ is used to subtract the noise $\epsilon_{\theta}(y_{t},t)$, resulting in a denoised image $Y_{t-1}$. This denoised image $Y_{t-1}$ is then input into the diffusion model to obtain the noise image for the next timestep. This process is repeated until $t=1$, at which point the denoised image is quite clear.
Sampling Process: In the inference process, which is also known as the sampling process, a new image $y_{0}$ can be generated from either Gaussian noise or a noisy image $y_{t}$ by iteratively sampling $y_{t-1}$ until $t=1$ according to the following expanded Eq. (9)
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},t)\right)+\hat{\beta}_{t}z$ (14)
where $z\sim\mathcal{N}(0,I)$, and $\hat{\beta}_{t}$ is usually approximated as $\beta_{t}$ in practical [15]. Such a sampling process is illustrated in Fig. 3.
Figure 4: The proposed taxonomy of diffusion model applications in RS.
II-B Conditional Diffusion Model
Similar to the development of GAN, the diffusion model was first proposed with unconditional generation, and then followed with the conditional generation [16]. Unconditional generation is often used to explore the upper limits of model capabilities, while conditional generation is more conducive to applications since it allows for the output to be controlled based on human wishes.
The first work to introduce the condition in a diffusion model is [44], which guides the generation of the diffusion model by adding a classifier to the well-trained diffusion model, so it is also called the Guided Diffusion Model. Although this method is less expensive to train, it also increases the inference cost by utilizing classification results to guide the sampling process of the diffusion model. More importantly, it has poor control over details and fails to produce satisfactory required images. As a result, the Google team [45] decided to adopt a straightforward idea to control the generated results by retraining the DDPM with conditions, and named it as Classifier-Free Guidance.
Formally, given conditional information $c$, the distribution of DDPM that need to be learned is changed as
$p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},c,t),\tilde{\beta}_{t}\mathbf{I})$ (15)
$\mu_{\theta}(x_{t},c,t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},c,t)\right)$ (16)
Correspondingly, the optimization objective (13) and sampling process (14) are modified into following forms:
${{\cal L}}_{con}(\theta)=\mathbb{E}_{x_{0},\epsilon,c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,c,t\right)\right\|^{2}\right]$ (17)
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},c,t)\right)+\hat{\beta}_{t}z$ (18)
Compared to the Guided Diffusion Model, this method is more widely used and is the basis of many attractive models (such as DALL-E2 [46], Imagen [47], Stable Diffusion [48] etc.), as well as the theoretical foundation of the conditional diffusion model in RS mentioned below.
III Applications of Diffusion Models in Remote Sensing
In this section, we will review and summarize existing related work, all of which involve the use of diffusion models in addressing RS image-related problems. To better organize our review, we categorize these papers according to their applications in RS and provide subdivisions for some common applications, as illustrated in Fig. 4. It is important to note that some applications may overlap with each other, but our categorization attempts to align with the core problems addressed by each paper.
III-A RS Image Generation
As one of the most impressive deep generative models, diffusion models are expected to synthesize realistic RS images from existing images or given textual descriptions to support the development of various RS applications. According to the data sources, these image generation methods can be mainly divided into two categories: text-to-image generation and image-to-image generation.
III-A1 Text-to-Image
Over the past two years, numerous text-to-image diffusion models have come out in computer vision [49, 47, 48, 16], especially the Stable Diffusion (SD) [48] model that has been widely adopted since its release [50, 51, 52]. However, its success mainly depends on training with billions of text-image pairs from the internet [53], which makes it difficult to extend to the field of RS since such vast and diverse RS datasets are not readily available. To address this problem, a straightforward idea is to produce trainable RS text-image pairs. Ou et al. [54] realized this idea with the help of pre-trained large models. Specifically, they first caption the existing RS images through a vision-language pre-training model to obtain initial textual prompts. Then, refining these prompts with human feedback and GPT-4 to improve semantic accuracy and suitability, successfully enabling the SD model to synthesize the required RS images. Instead of directly supplementing text prompts, Khanna et al. [55] used various numerical information related to satellite images, such as geolocation and sampling time as new prompts of the SD model, which effectively enriches the SD model’s input and enhances its ability to generate high-quality satellite images. Furthermore, Tang et al. [56] refined the generation process with SD model by incorporating RS image-related features as control conditions. They treated textual descriptions and numerical information as global control information, and used the depth map, segmentation mask, object boundaries, and other result images obtained through a series of pre-trained networks as local control information. By flexibly selecting control conditions, this approach achieves effective integration of multiple control information, expanding the RS image generation space.
Despite the significant progress made by the improved SD model on optical RS images, researchers recently encountered new challenges when adapting it to other modal RS images [57]. For example, directly using SAR images to fine-tune the SD model will degrade the model’s representational ability, resulting in the failure to generate satisfactory SAR images. This is because there are significant differences in capture perspectives and data modalities between SAR and natural images. In view of this, Tian et al. [57] proposed to fine-tune the SD model with optical RS images before using SAR images, so as to transition the model from regular view to the bird’s-eye view. Meanwhile, they suggested training the SD model’s Low-Rank Adaptation network [58], rather than the whole model, thus ensuring that semantic knowledge learned from natural images can be successfully transferred to the learning process of SAR images.
Apart from fine-tuning the SD model, researchers in the RS community have attempted to design a new architecture for RS text-to-image generation with diffusion model [59]. The proposed pipeline consists of two cascaded diffusion models, where the first one is designed to generate low-resolution satellite images from text prompts, and the second one is to increase the resolution of the generated images based on the text descriptions. The benefits of this two-stage generation approach are twofold. On the one hand, separating the generation of low and high-resolution images is advantageous for capturing scene information from global to local perspectives. On the other hand, the low-resolution image generation stage can alleviate the computational burden of generating high-resolution satellite images directly from text descriptions, which is more feasible for practical deployment.
III-A2 Image-to-Image
Compared to the text-to-image generation, image-to-image generation is more popular in the field of RS. In this task, diffusion models are guided by existing images to generate new ones. The guiding images can take various forms. Some researchers prefer using masks, such as maps [60], class labels [61], and semantic layouts [62, 63, 64, 65], as the guiding images for the diffusion models. Although these images only contain some specific information, they still present excellent performance in the RS image generation. For example, [60] successfully produces realistic satellite images by training the ControlNet model [66] with maps, even historical maps. [61] effectively addresses the issue of insufficient samples and unbalanced classes captured from actual battlefield environments by using class labels as conditional constraints. [63] achieves the generation of RS image-annotation pairs by implementing a two-phase training process on the SD model, addressing the expensive high quality annotation problem. One of the most noteworthy works is [65], which not only generates high-quality RS images with the guides of semantic masks, but also addresses the inherent problem of diffusion models requiring long training time for the model convergence. Specifically, this work introduces a lightweight diffusion model obtained through a customized distillation process, which ensures the quality of image generation via a multi-frequency extraction module and achieves rapid convergence by adjusting the image size at different stages of the diffusion process.
Figure 5: Overview of different diffusion model-based methods for RS image generation. Note that also combinations of different condition inputs are possible.
However, the masks used in these methods are essentially annotation labels that require expert knowledge and manual labeling, making them costly to acquire. Given the difficulty in acquiring these masks, some works have explored to use multi-modal RS images as the guide images of diffusion models [67, 68, 69]. Different modal RS images have their own strengths and weaknesses. For example, optical RS images are highly visualized and can intuitively reflect surface information, but they are limited by weather conditions and capture time, and are easily obscured by clouds. Conversely, SAR images can be captured in all weather conditions and penetrate clouds and fog, but their imaging process is complex and usually require experts to interpret. In view of these, Bai et al. [67] adopted SAR images as the guide images for a diffusion model to generate optical RS images, which gains higher clarity and better structural consistency than those generated by GAN models. Similarly, hyperspectral images (HSIs) can provide richer spectral-spatial information than multispectral images (e.g., RGB images), even though both belong to optical RS images. However, the acquisition cost of HSIs is much higher than that of multispectral images. Therefore, researchers would like to generate HSIs with the help of easy-obtainable multispectral RS images [68, 69]. Unfortunately, using diffusion model to generate HSIs requires matching the input noise dimensionality with the spectral bands of the HSI, resulting in an excessively large noise sampling space that hampers the model’s convergence. To address this issue, Zhang et al. [68] proposed a spectral folding technology to convert the input HSI into a pseudo-color image before training the diffusion model. Liu et al. [69] used conditional vector quantized generative adversarial network (VQGAN) [70] to obtain a latent code space of HSIs, and performed the training and sampling processes of diffusion model within this space.
Overall, the above image-to-image generation methods are based on the conditional diffusion model [45], using the guide image as the condition input to the conditional DDPM, which essentially generates target RS images from noise rather than images directly. To realize the true image-to-image translation, Wang et al. [71] employed a straightforward idea: input Inverse Synthetic Aperture Radar (ISAR) images [72] to the diffusion model during the training phase to force the model to learn the distribution of ISAR images, and then input optical RS images in the testing phase to generate new ISAR images. Seo et al. [73] proposed to generate optical RS images from SAR images by sampling noise from the target images rather than using Gaussian noise, efficiently ensuring the consistency of the distribution between the generated image and the target image without using the conditional diffusion model. More recently, Li et al. [74] the diffusion model for generating 3D urban scenes from satellite images. They utilized a 3D diffusion model with sparse convolutions to generates texture and colors for the foreground point cloud, comprising buildings and roads, and employed a 2D diffusion model to synthesize the background sky, which enabled the direct generation of 3D scenes solely from satellite imagery, demonstrating a novel application of diffusion models in the RS community. Apart from generating images from one modality to another, researchers have proposed that large satellite images can be generated from patch-level images within the same modality [75]. Specifically, they first adopted self-supervised learning to extract feature embeddings of the input patches, and then used these embeddings as condition input to guide the diffusion model’s learning. Notably, they retained the spatial arrangement of each patch in the original image, making it possible to assemble the generated patches into a large and coherent image.
III-B RS Image Enhancement
III-B1 Super-Resolution
Remote sensing super-resolution (RSSR) aims to reconstruct high-resolution (HR) RS images with more details from low-resolution (LR) RS images [76], which are degraded by imaging equipment, weather condition, or downsampling. Compared to natural images, RS images suffer from much more details loss, making it more challenging to reconstruct the HR images [77]. Therefore, using a more powerful generative model, such as the diffusion model, to complete the RSSR task has received widespread attention in the RS community. Fig. 6 compares the workflow of previous deep learning-based and diffusion model-based RSSR methods. Unlike the previous deep learning-based methods, which attempt to find a suitable mapping function to fuse the detail information in just one step, diffusion model-based methods integrate the information from LR image and guided image into each step of the diffusion process, which facilitates the better a better fusion of different image information.
Figure 6: Comparison of previous deep learning-based methods and diffusion model-based methods for the RSSR task. (a) The workflow of previous deep learning-based methods, where L, G, H indicate the LRRS image, guided image and HRRS image. (b) The workflow of diffusion model-based methods, where $\textbf{X}_{t}$ represents the diffused HR image at t and G represents the condition input (such as the LR image and its features) for the diffusion model. The figure originally shown in [78]
Multispectral images
Given that RS images always contain small and dense targets, Liu et al. [77] proposed a diffusion model with a detail supplement mechanism for the RSSR task, which requires two-step training to realize. Specifically, the first training aims to improve the model’s capability for reconstructing small objects through randomly masking HR images, and the second training is to complete the super-resolution task by utilizing a conditional diffusion model with LR images as condition input. Although superior performance is achieved, the dual training process in this method is complex and time-consuming. To simplify the training process, Han et al. [79] leveraged Transformer [80] and CNN to extract global features and local features from LR images, respectively, and used the fused feature images to guide the diffusion model generate HR images. In this way, the function of two-step training in [77] is successfully realized in one training. Similarly, Xiao et al. [81] extracted rich prior knowledge from the original LR images by using stacked residual channel attention blocks [82] to guide the optimization of the diffusion model. Furthermore, An et al. [83] departed from the commonly used U-Net architecture in diffusion models, implemented an encoder-decoder architecture through parameter-free approaches, and adopted denoising diffusion implicit models (DDIM) [84] to accelerate sampling process, which significantly improves the efficiency of generating HR images, and is more suitable for diverse RS scenarios.
The above RSSR methods are designed on the assumption that LR images are generated from a fixed degradation model, such as downsampling. However, the blurring in real remote sensing images are complex and varied, which can be modeled as many different degradation models. In view of this, Xu et al. [85] proposed to solve this problem with two diffusion models, where the first one is trained as a degradation kernel predictor, so that the predicted degradation kernel and LR image can be used together as conditions in the second diffusion model to generate the HR images. Feng et al. [86] achieved the learning of degradation kernel and the reconstruction of HR images within a single diffusion model through the use of the kernel gradient descent module and kernel proximal mapping module.
HSIs
Although HSIs possess high spectral resolution, their spatial resolution is relatively low [87], which may limit the performance of various applications based on HSIs. To obtain high spatial resolution HSIs, there are two categories methods: pansharpening [88] and multispectral and hyperspectral image fusion [89, 90, 91, 92]. As the name suggests, the former enhances the HSI by injecting the detail information from panchromatic (PAN) images, while the latter leverages multispectral images to help the HSI learn spatial details. For example, Shi et al. [91] used the concatenated image of the multispectral and HSI as the condition input for the diffusion model, enabling the model to capture useful information from both image modalities to generate HSIs with high spatial resolution.
Compared to using multispectral images as the detail guides, researchers are increasingly dedicated to achieve HSI super-resolution with PAN images [93, 78, 94, 95, 96, 97, 98]. Instead of crudely concatenating the PAN and HSI images directly, Meng et al. [93] proposed a Modal Intercalibration Module to enhance and extract features from both images, where the enhanced features is used as condition input to the diffusion model. Cao et al. [78] believed that the unique information in different image modalities should not be blended for processing. Thus, they designed two conditional modulation modules to extract coarse-grained style information and fine-grained frequency information respectively as the condition inputs. Still for fully utilizing the unique information of different modalities, Li et al. [95] proposed a dual conditional diffusion models-based pansharpening network, which takes HSI and PAN as independent condition inputs to learn the spectral features and spatial texture respectively. Notably, the training and sampling of the proposed diffusion model are performed in a low-dimensional latent space constructed by an auto-encoder network, which not only reduces the computational cost but also maintains the model’s generalization ability. [96] also performs the sampling process in a low-dimensional space. Based on the assumption that the HRHSI can be decomposed into the product of two low-rank tensors, this method first computes one of the low-rank tensors with the LRHSI. Then, this tensor is taken as the condition along with the LRHSI and PAN, input into a pre-trained RS diffusion model [99] for another low-rank tensor. This method, unlike the above-mentioned methods, is a completely unsupervised deep learning method that does not require the involvement of HRHSI in any process, providing feasibility for its application in practice.
Figure 7: Overview of diffusion model-based methods for RS image denoising. Since multispectral and hyperspectral images typically exhibit additive noise, diffusion models can be directly employed for denoising. In contrast, SAR images are corrupted by the multiplicative noise. Consequently, there are three approaches to address this problem: directly utilizing the diffusion model, transforming the multiplicative noise into additive noise, as well as transforming the diffusion model to be fit for removing multiplicative noise.
III-B2 Cloud Removal
In many cases, optical RS images will be partially obscured by clouds since the view of imaging equipment is limited. Cloud removal, in essence, is to reconstruct the areas corrupted by clouds, which means to generate content that is consistent with the surrounding environment to fill the missing areas of the image [100]. Therefore, it is suitable to use diffusion model for this task, since it can better control the generated content.
The control condition for cloudy image reconstruction comes in various forms. For example, Czerkawski et al. [101] adopted text prompts and edge information [102] as the guiding conditions, together with the input cloudy image, cloud mask, and diffused cloud-free image to control the generation process of SD [48]. Jing et al. [103] input both SAR image and cloudy optical RS image into the diffusion model for feature extraction, effectively enhancing the cloud removal results with the help of cloud-unaffected SAR image. Zou et al. [104] first extracted features from the cloudy image and noise level, and then inputs the extracted spatial and temporal features into the diffusion model as control conditions. Instead of using more control conditions, they trained the model in an supervised manner with cloud-free images to further improve the quality of the reconstructed images. Moreover, Zhao et al. [105] integrated different images from multi-modalities and multi-time into a sequence input, and utilized a two-branch diffusion model to extract scene content from the optical RS and SAR images, respectively. Unlike the previous works, this method does not require the participation of cloud-free images and can handle image sequences of any length, offering greater flexibility and practical value.
III-B3 Denoising
Due to the inherent constraints of imaging technology and environmental conditions, RS images are always accompanied by various noise [106]. In essence, the learning process of diffusion model is equal to the denoising process [15], which makes it possible to achieve superior denoising performance in the context of RS. Different modalities of RS images confront different noise challenges. For example, optical RS images are more susceptible to blurring caused by atmospheric scattering and absorption [107, 108], as well as the noise from changing lighting conditions [109]. HSIs need to consider the uneven distribution of noise over the spectral dimensions [110] and the correlation of noise between different bands [111, 112]. Fortunately, these noises can be effectively removed through diffusion models. Huang et al. [107] proposed to crop the noisy RS image into small regions and rearrange them in a cyclic shift manner before fed into the diffusion model, so as to achieve finer local denoising as well as artifact elimination. He et al. [111] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process, instead of a pure noise, to avoid the destruction of the inherent effective information in the HSI. Moreover, Yu et al. [113] simulated the harsh imaging conditions of RS satellites by adding various attack disturbances to input images, enhancing the diffusion model’s ability to counteract the system noise.
SAR images, as another modality of RS images, are usually contaminated by a multiplicative noise, speckle [114]. Unlike additive noise, the degradation caused by speckle varies across different areas in one image. Therefore, speckle noise significantly affects the disparity and interdependence between pixels, causing severe damage to the image. To eliminate this particular noise, Perera et al. [115] proposed to use the speckled SAR image, along with the Gaussian noised SAR image that conforms to the standard diffusion model for denoising training. However, this method of introducing the synthetic noise images may not be able to accurately and reasonably simulate the actual SAR images, causing suboptimal denoising performance. Thus, Xiao et al. [116] proposed to transform the multiplicative noise in SAR into additive noise by log function, enabling the transformed SAR images to match the standard diffusion model for independent training. Further advancing this methodology, Guha et al. [117] integrated the log operation into the derivation process of the diffusion model and obtained a one-step denoising network that can directly address the multiplicative noise.
Figure 8: Overview of diffusion model-based methods for landcover classification. For multispectral images, the ground-truth map is typically used as the diffused image, with the original RS image serving as the condition. The dashed line indicates that the classification results predicted by the network can also be used as the diffused image for the next input. As for HSIs, the prevailing approaches involve using the diffusion model as a feature extractor, after which the extracted features are fed into a classifier for classification.
III-C RS Image Interpretation
III-C1 Landcover Classification
As one of the most common applications in the RS community, landcover classification aims to assign each pixel to a specific class, such as buildings or grass, to obtain useful landcover information. However, the RS images always contain diverse and complex scenes, increasing the difficulty of accurate classification. Given that diffusion model can learn and simulate complex data distributions better than other deep learning models, researchers are trying to use it for the RS landcover classification.
The first application of the diffusion model to this task is presented in [118], based on the conditional diffusion model [44]. It takes the manually annotated ground-truth map and the original RS image as the diffused image and condition respectively, and decouples the commonly used U-Net architecture by adding two separate encoders for extracting features from both diffused and guiding images. Besides, it averages the results of multiple sampling for the final classification result to improve the stability and overall accuracy. This method was later tested by Ayala et al. [119] on wider RS datasets, validating the effectiveness and development potential of the diffusion model in the RS landcover classification. Instead of using the ground-truth map as the diffused image, Kolbeinsson et al. [120] diffused the classification prediction result of the network from the previous step and input it along with the conditioning RS image into the diffusion model for next classification prediction. Notably, the parameters of their diffusion model are optimized not only through the MSE of predicted noise but also by minimizing the difference between the prediction result and ground-truth map at each step.
Compared to multispectral images, HSIs exhibit a more complex data distribution, posing a greater challenge for the application of conventional deep learning models in landcover classification. Fortunately, the diffusion model can better capture the spectral-spatial joint features of HSIs, facilitating the improvement of classification accuracy [121]. Based on this fact, Zhou et al. [122] constructed a timestep-wise feature bank by utilizing the temporal information of the diffusion model, and proposed a dynamic fusion module to integrate spectral-spatial features with temporal features, making it possible to obtain sufficient image information before the classification. Li et al. [123] designed a dual-branch diffusion model for feature extraction from HSI and LiDAR images separately, achieving information complementarity between different modal RS images and enhancing the distinguishability among pixels, which also demonstrated superior performance in a multi-client RS task [124]. Qu et al. [125] further set the encoder of the dual-branch diffusion model to operate in parameter-sharing mode to ensure the extraction of shared features from multimodal RS images. In addition, Chen et al. [126] utilized the diffusion model to assist deep subspace construction, achieving excellent HSI classification performance in an unsupervised manner. Different from the aforementioned methods, Ma et al. [127] used the diffusion model to directly classify the HSI pixels into background and anomaly targets, rather than as a feature extractor. Specifically, they adopted a diffusion model to learn the background distribution based on the fact that the distribution of the background obeys a mixed Gaussian. Thus, the background is removed as noise during the inference process, effectively retaining the anomalous pixels of interest.
Figure 9: Comparison of (a) two-step and (b) end-to-end diffusion model-based CD methods. The two-step methods separate the training process of diffusion model and CD head, where the diffusion models (in yellow) are pre-trained by millions of RS images and only the CD head is optimized by the CD dataset. In contrast, the end-to-end method enables the joint optimization of diffusion models and CD module. The figure originally shown in [128]
III-C2 Change Detection
RS change detection (CD) aims to identify the differences between two images taken at different times of the same area [129], thereby providing support for environmental monitoring, and natural disaster assessment. Considering the exceptional performance of diffusion models in handling image details, such as textures and edges, which are crucial for distinguishing changes, some researchers have explored diffusion model-based CD methods.
Bandara et al. [99] employed a pre-trained diffusion model to extract multi-scale features of different temporal images for the CD module training, where the pre-training was accomplished by millions of free and unlabeled RS images to capture the key semantic. Tian et al. [130] integrated a diffusion model into the contrastive learning framework [131] to capture fine-grained information in the RS images, successfully extracting features with clearer boundaries and richer texture details for the CD task. Additionally, Zhang et al. [132] adopted the Transformer [133] as the backbone for the diffusion model to extract spectral-spatial features from HSIs captured at different times. In essence, all of the above methods use the diffusion model as a feature extractor trained separately from the CD task, which results in the generated features not being fully suitable for CD and overlooks the potential benefits of gradual learning and controllability provided by the diffusion model. To address these issues, Wen et al. [128] and Jia et al. [134] proposed an end-to-end diffusion model for CD, where the diffused ground-truth map is input into the network along with the pre- and post-change images, and the difference between two temporal images is used as the condition to guide the direction of detection.
III-C3 Climate Prediction
Climate prediction is another application for diffusion model in the field of RS, which is a complex systematic project that requires the integration of multiple variables, including cloud amount, cyclone distribution, and water vapor. Therefore, its solution often consists of multiple diffusion models, enabling fine-grained, stepwise processing of these high-dimensional data. For example, Nath et al. [135] cascaded three independently trained diffusion models to generate future satellite imagery, increase the resolution, and predict precipitation. Hatanaka et al. [136] utilized two cascaded score-based diffusion models to generate high-resolution cloud cover images from coarse-resolution atmospheric variables. In addition, Leinonen et al. [137] addressed the issue of high computational costs by adopting the latent diffusion model concept proposed in [48], running the diffusion process within a latent variable space mapped by a 3D VAE network. These methods all demonstrate that diffusion models can better capture the complex spatio-temporal relationships and become a powerful tool in climate prediction.
III-C4 Miscellaneous Tasks
Apart from the classic RS interpretation tasks reviewed above, there are some other tasks that may not fall into the above categories [138, 139, 140].
One of such tasks is height estimation, which aims at providing pixel-wise height information of surface features (e.g., buildings, trees, terrain, etc.) to generate 3D models of surface scenes. Diffusion model has been reported to be a promising solution for surface estimation [138]. Unlike traditional methods that require multi-view geospatial imagery or LiDAR point clouds, it produces accurate height estimates only with single-view optical RS images.
The other task is object detection, which uses bounding boxes to locate the instances of a certain class (such as plane, vehicle, or ship) in RS images. One of the most challenging issues in RS object detection is the lack of sufficient training data [141]. This scarcity is due to the long-distance photography of RS images, which often results in the object of interest being small and sparsely distributed across different regions in the images. Thus, augmenting the object of interest with the diffusion model has become an effective solution [139]. Specifically, it works by training the SD model with object patches cropped from the available object detection training set, where the patches are 10 pixels larger than the corresponding ground-truth boxes to ensure that sufficient context is captured for seamlessly merging with the background. Compared with other data augmentation methods, this method can extract the precise coordinates of the synthesized objects, and effectively mitigates the long-tailed distribution of target categories.
Additionally, diffusion model also shows superior performance over CNNs in the task of anomaly detection in satellite videos (such as wildfire detection) [140]. The past frames serve as the condition input for the diffusion model, enabling it to learn the data distribution of normal frames and generate high-quality data that closely resemble real images. Consequently, when an anomalous frame is input, the model outputs a significantly higher anomaly score. This means that the diffusion model can detect small wildfires promptly, preventing widespread fire outbreaks, which is as opposed to CNN-based methods that usually require the fire to reach a certain visual extent to be effective.
Figure 10: Comparison of different cloud removal methods on the Sen2-MTC-Old dataset: (a) Cloudy Image T1. (b) Cloudy Image T2. (c) Cloudy Image T3. (d) Ground-Truth. (e) STNet [142]. (f) DSen2-CR [143]. (g) PMAA [144]. (h) UnCRtainTS [145]. (i) DDPM-CR [103]. (j) DiffCR [104]. The visual results are retrieved from [104].
III-D Experimental Evaluation
To effectively illustrate the superiority of diffusion models in processing RS images, we take the experimental results of cloud removal, landcover classification, and change detection as examples in this section to evaluate the performance of the diffusion model and other existing techniques through visual results and quantitative indicators.
TABLE I: Quantitative Comparison of Different Cloud Removal Methods on The Sen2-MTC-Old Dataset
Methods Metrics
PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$
STNet [142] 26.321 0.834 146.057 0.438
DSen2-CR [143] 26.967 0.855 123.382 0.330
PMAA [144] 27.377 0.861 120.393 0.367
UnCRtainTS [145] 26.417 0.837 130.875 0.400
DDPM-CR [103] 27.060 0.854 110.919 0.320
DiffCR [104] 29.112 0.886 89.845 0.258
*The results are retrieved from [104].
Figure 11: Comparison of different HSI classification methods on the Salinas dataset: (a) Pseudo-Color Image. (b) Ground-Truth. (c) SF [146]. (d) miniGCN [147]. (e) SSFTT [148]. (f) DMVL [149]. (g) SSGRN [150]. (h) SpectralDiff [121]. The visual results are retrieved from [121].
Figure 12: Comparison of different change detection methods on the LEVIR dataset: (a) Pre-change Image. (b) Post-change Image. (c) Ground-Truth. (d) FC-SD [151]. (e) STANet [152]. (f) SNUNet [153]. (g) BIT [154]. (h) ChangeFormer [155]. (i) DDPM-CD [99].
Fig. 10 displays the visual results of six different cloud removal methods on multi-temporal optical satellite (Sentinel-2) images [156], where the first three cloudy images are taken at different times from the same location. As shown in Fig. 10, diffusion model-based methods DDPM-CR [103] and DiffCR [104] have successfully removed clouds without leaving excessive artifacts, restoring the RS image with detailed information. This observation is also confirmed by the quantitative indicators in Table I. Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [157], Learned Perceptual Image Patch Similarity (LPIPS) [158] and Frechet Inception Distance (FID) [159] are used in Table I to evaluate the quality of cloud-free images generated by the comparison methods. As shown in Table I, DiffCR achieves the best performance on all four indicators, and DDPM-CR, another diffusion model-based method, also ranked second on FID and LPIPS indicators, which demonstrates that using diffusion models for cloud removal is highly competitive.
TABLE II: Quantitative Comparison of Different HSI Classification Methods on The Salinas Dataset
Methods Metrics
OA (%) AA (%) $\kappa$ (%)
SF [146] 88.248 93.262 86.973
miniGCN [147] 88.181 94.297 86.823
SSFTT [148] 95.789 98.272 95.322
DMVL [149] 97.005 95.853 96.668
SSGRN [150] 96.539 96.354 96.144
SpectralDiff [121] 98.971 99.465 98.854
*The results are retrieved from [121].
TABLE III: Quantitative Comparison of Different Change Detection Methods on The LEVIR Dataset
Methods Metrics
F1(%) IoU(%) OA(%)
FC-SD [151] 86.31 75.92 98.67
STANet [152] 87.26 77.40 98.66
SNUNet [153] 88.16 78.83 98.82
BIT [154] 89.31 80.68 98.92
ChangeFormer [155] 90.40 82.48 99.04
DDPM-CD [99] 90.91 83.35 99.09
*The results are retrieved from [99].
As illustrated in Fig. 11, the diffusion model-based method SpectralDiff [121] has significantly better classification results than SF [146], miniGCN [147], and SSFTT [148] on the hyperspectral dataset Salinas. Although DMVL [149] and SSGRN [150] show comparable performance to SpectralDiff across most classes, they are not as accurate as SpectralDiff in assigning pixels at the boundaries of different classes. In addition to the visualized results, Table II lists the quantitative results of these methods on overall accuracy (OA), average accuracy (AA), and Kappa coefficient, where SpectralDiff ranks first.
As for the change detection task, we selected five comparised methods based on different deep learning models along with a diffusion model-based method, DDPM-CD [99], for experimentation on the LEVIR dataset [160]. The corresponding visual results and the values of evaluation indicators, i.e., F1 score (F1), overall accuracy (OA), and intersection over union (IoU), are presented in Fig. 12 and Table III, respectively. It is evident from both qualitative analysis and quantitative comparisons that the change detection method based on the diffusion model is significantly superior to the others.
IV Discussions and Future Directions for RS Diffusion Models
As discussed in the previous sections, diffusion models are rapidly evolving in the RS community, presenting great potential from generating RS images to enhancing the image quality, and further to recognition and detection. In fact, the research on diffusion models in RS is still at an early stage, with many tasks to be explored and further improvements to be achieved. In the following sections, we will discuss the possible future research directions from two aspects: extended applications and model deployment.
Figure 13: Frequency of diffusion models in different RS applications.
IV-A Extended Applications
IV-A1 For Specific RS Tasks
As shown in Fig. 13, diffusion models are most frequently used in the RS image generation task, followed by the task of reconstructing high-resolution RS images (i.e., the super-resolution task), which accounts for 21% of all reviewed papers. Comparatively, the use of diffusion models in advanced and complex interpretation tasks is quite less. Especially for object detection, an important and common RS task, there is only one paper related to the diffusion model [139], which implies a huge gap in the research of diffusion models for this task. Recently, Chen et al. [161] proposed a new object detection network named DiffusionDet, which finds the correct positions of objects by gradually denoising from random noise boxes to object boxes. This method ingeniously combines the object detection task with diffusion model, and achieves favorable performance on the COCO dataset [162]. Although the performance of this method on small and sparse RS targets is still unclear, its emergence has paved an enlightening new path for researchers in the RS field to further explore in the object detection task. Nevertheless, Fig. 1 presents that the research on diffusion models for RS image generation has decreased since 2024, while research on their application in RS image interpretation tasks, especially in the change detection task, has significantly increased. This suggests that researchers have realized the limitations of diffusion models in the application to RS. Therefore, developing more effective RS diffusion models for image interpretation tasks is an important research direction for the future.
While diffusion models are also commonly applied in landcover classification and change detection, there exists a significant limitation. In these methods, diffusion models mainly serve as feature extractors for the input image [121, 123, 122, 124, 132, 99, 130], requiring an additional classifier/detector to execute the specific task. Such separation of feature extraction and detection is not straightforward and usually needs two-step training, which is likely to fall into a suboptimal solution due to the extracted features not being fully suitable for the specific task. Therefore, future research could focus on developing RS task-specific diffusion models, by incorporating task-related prior in the model design or designing end-to-end models to improve the performance of diffusion models on specific RS tasks.
It is worth noting that the vast majority of existing RS diffusion models are based on the U-Net architecture, with only a few works incorporating the Transformer architecture or its attention mechanism [79, 132, 121]. Among the most innovative is [132], which employs a completely Transformer-based diffusion model, U-ViT [133], for change detection. Nevertheless, the U-ViT consists of long skip residual connections, aligned with those of U-Net. In contrast, DiT [163], another diffusion model entirely based on the Transformer architecture, set the residual connections within each block, allowing the attention layer to perform global convolution and information extraction at a finer-grained scale, achieving state-of-the-art results in both image and video generation [163, 164, 165, 166]. Thus, it can be seen that the diffusion model based on transformer architecture hold great potential for development, warranting further exploration by researchers in the field of RS.
TABLE IV: Categorization of Diffusion Models in RS Based on Image Modalities
Modality Application Correlation paper
Multispectral Generation
[55][59][54][65][61][62][75]
[60][67][73][64][56][63]
Cloud Removal [100][101][103][104][105]
Super-Resolution [77][79][81][83][85][86]
Hyperspectral Super-Resolution
[78][90][91][92][93][94][95]
[96][97][98]
Classification
[121][122][123][124][125]
[126][127]
Denoising [111][112][110]
SAR Denoising [115][116][117]
Generation [57][71][167]
IV-A2 For Multi-Modal RS Images
Unlike natural images, RS images are captured by different types of sensors, encompassing multiple modalities. Table IV lists the top three most common applications across different modal RS images. Obviously, most of the diffusion model-based methods are developed for multispectral images. For HSI images, the applications of diffusion models are focused on super-resolution, especially pansharpening [93, 78, 94, 95, 96], and classification tasks. As for SAR images, diffusion models only appear in the process of generation and denoising. However, the spectral signatures of HSI and the robustness of SAR images play crucial roles in the recognition and detection tasks [168, 169, 170, 171, 172]. Therefore, exploring how to effectively apply diffusion models to multi-modal RS images is a necessary future research direction. In this way, the unique information of different modalities can be leveraged to improve the accuracy of RS image analysis.
In addition, LiDAR data, as an important type of RS data that can provide surface height information and ground structure details, is seldom used in the existing RS diffusion models. Only Li et al. [123] introduced LiDAR images as auxiliary information when using diffusion models for hyperspectral classification. In fact, LiDAR images can not only be used as supplement to other modalities [173, 174], but also generate continuous 3D terrain models for topographic and geomorphological analysis [175], vegetation detection [176, 177], and urban planning [178, 179]. Recently, diffusion models have demonstrated satisfactory performance in 3D point cloud generation [180, 181, 19]. Such technological advancements may be borrowed to the RS LiDAR data, thereby filling the gap of LiDAR data in various RS tasks.
IV-A3 For Realistic RS Images
Although many diffusion model-based RS image generation methods have been developed, they often overlook some special features of RS images, resulting in noticeable gaps between the synthesized image and real images. For example, diffusion model-based HSI generation methods usually need to compress the spectral dimension [68, 69], which neglects the details of spectral curves and hinders the generation of accurate spectra. SAR images are complex in nature, comprising both amplitude and phase terms [182]. However, the generation methods developed so far have mainly focused on the image amplitude, ignoring the phase information. Therefore, future diffusion model-based generation methods should take these aspects into account to obtain more realistic RS images.
Another aspect that is often overlooked is the size of real RS images, which tend to be quite large (e.g., Gaofen-2 images are 29,200 $\times$ 27,620 pixels) [183, 184]. Existing diffusion model-based generation methods are primarily designed for patch-level images with 256 $\times$ 256 pixels, which means that a large-scale RS image can only be obtained by stitching multiple patch images [75]. From the visual perspective, this approach is suboptimal since it is difficult to ensure the continuity of scenes and easy to leave traces at the joints. Thus, how to obtain realistic and reasonable large-scale RS images with diffusion models is a worthy research direction for further exploration. Notably, the increase in image size will inevitably bring computational and storage burdens. Accordingly, how to generate large-scale RS images under limited resources also needs careful consideration.
IV-A4 For General RS Model
Nowadays, more and more researchers are devoting themselves to developing general intelligent models, which can provide more accessible and high-performing solutions to help both industry professionals and interested non-professionals [185, 186, 187, 188, 46]. For example, ChatGPT [187] has greatly simplified the process of collecting and summarizing information, while DALL-E [188] has facilitated the rapid transformation of artistic ideas into practical examples. In the field of RS, some researchers are also attempting to develop universal multi-modal large models [189, 190, 191, 192]. However, these studies are still in the embryonic stage and have a lot of room for development. Fortunately, diffusion models have shown superior performance on various RS applications, as presented before. Therefore, a promising research direction is to construct a general RS intelligent model based on the diffusion model, which can span different modalities of RS images and accomplish multiple earth observation tasks.
IV-B Model Deployment
It is widely acknowledged that diffusion models require a substantial number of iterations to generate high-quality samples, which is a noticeable drawback of this technology. Moreover, these models often encompass a substantial number of parameters, necessitating deployment on devices with powerful neural computing units [193, 194]. However, the resources on satellites are extremely limited and lack the computational power for such requirements. What’s more serious is that the deep-space environment is severe, often subject to extreme climate and illumination changes, which puts higher demands on the stability and reliability of the model. Therefore, how to deploy diffusion models on resource-limited satellites within harsh environments for real-time processing is a very meaningful research direction.
IV-B1 Accelerate Processing
Recently, some accelerated sampling approaches for diffusion models have been proposed [195, 196, 84, 197, 198], which successfully reduce the necessary sampling steps from several hundred to dozens, or even a few or a single step [199, 200, 201]. However, Tuel et al. [167] found that these accelerated sampling methods, such as DDIM [84] or DPM-solver [197], did not perform well on SAR images. This observation suggests that the significant differences between RS images and natural images make these acceleration methods, originally designed for natural images, inapplicable to RS imagery. More recently, Kodaira et al. [202] proposed StreamDiffusion, which achieves the speed of generating images up to 91.07fps on a 4090GPU through pipelined batch processing. Unfortunately, they did not evaluate the proposed model on RS images and overlooked that some of the acceleration techniques are not suitable for devices with limited computing resources. It can be seen that developing acceleration methods for RS diffusion models is still an open problem.
Another common optimization strategy is to design lightweight diffusion models, which can be achieved by distillation techniques [65, 79, 203, 204] or changing the network structure [83, 81, 205, 206]. Nevertheless, the more prevalent solution in RS is to place the diffusion model in a lower-dimensional latent space through data compression [95, 96, 69, 137] thereby shrinking the sampling space and reducing computational overhead. However, the performance of such methods is limited by the quality of the constructed latent space [207]. If the latent space is incapable of extracting useful semantic information from RS imagery, or critical information is lost during the dimensionality reduction process, the final generated RS images will be adversely affected. Therefore, it is necessary to explore better lightweight structure design methods than compressing the data space.
IV-B2 Improve Stability and Reliability
As for the deployment challenges posed by the harsh environment, Yu et al. [113] proposed a diffusion model-based adversarial defense approach to protect deep neural networks from a variety of unknown adversarial attacks, effectively improving the robustness of the model. Similar research on employing diffusion models to tackle the complex environments in RS should be further explored, especially to design specific diffusion models for different environmental disturbances or different sensors. Such efforts would contribute to the long-term operational stability of the deployed models and hold great practical value.
In addition, with the rapid development of Global Observation System (GOS), the distributed learning of intelligent models over multiple satellites has gradually become one of mainstream directions in RS [208, 209, 210, 211]. More recently, Li et al. [212] reported that the speed of generating images with diffusion models can be effectively improved under this parallel learning mode, further confirming prospects of distributed diffusion models in RS. However, when deploying diffusion models in such multi-client distributed scenarios, ensuring the security of RS data is an essential issue [213]. Consequently, developing robust and trustworthy diffusion models has become an urgent necessity in the RS community.
V Conclusion
In summary, the emergence of diffusion models has created a new era of intelligent RS image processing. Compared to other deep generative models, diffusion models are robust to the inherent noise in RS images, better adapt to their variability and complexity, and offer a more stable training process. Hence, applying diffusion models to various RS tasks has become an inevitable trend. For this reason, this paper first introduces the theoretical background of diffusion models to help understand how the diffusion model works for RS tasks. Then, it reviews and summarizes studies on the use of diffusion models in processing RS images, including image generation, super-resolution, cloud removal, denoising, and a series of interpretation tasks such as landcover classification, change detection, and climate prediction. Moreover, the paper takes cloud removal, landcover classification, and change detection as examples to demonstrate the superiority of diffusion models in various RS image processing tasks through visual results and quantitative indicators. Finally, the paper discusses the limitations of the existing diffusion models in RS and highlights that further exploration could be carried out on the extended applications and model deployment. We hope this paper can provide a valuable reference for researchers in related fields to stimulate more innovative studies to break the performance bottleneck of existing methods or to promote the development of diffusion models for more RS applications.

### v2

@@
 In summary, diffusion models possess great development potential in the field of RS.
-Therefore, it is necessary to review and summarize existing diffusion model-based RS papers to help researchers gain a comprehensive understanding of current research status, and identify the gaps in the application of diffusion models in RS, thereby promoting further development in this field.
+Therefore, it is necessary to review and summarize existing diffusion model-based RS papers to help researchers gain a comprehensive understanding of the current research status, and identify the gaps in the application of diffusion models in RS, thereby promoting further development in this field.
 The remainder of this paper is organized as follows.
@@
 II Theoretical Background of Diffusion Models
-Diffusion models, also known as diffusion probabilistic models, are a family of deep generative models.
-In general, generative model is to convert a random distribution (i.e., noise) into a “probability distributions” that is the same as the distribution of the observed dataset, and obtain the desired outcomes by sampling from this “probability distribution”.
-Obviously, it is quite difficult to obtain the target distribution directly, but disrupting a regular distribution into random noise is straightforward, and can be achieved by continuously adding Gaussian noise, as illustrated in Fig.
+Diffusion models, also known as diffusion probabilistic models, constitute a family of deep generative models.
+In essence, a generative model aims to transform a random distribution (i.e., noise) into a ’probability distribution’ that matches the distribution of the observed dataset, thereby producing desired outcomes through sampling from this ’probability distribution’.
+Obtaining the target distribution directly is evidently challenging.
+However, disrupting a regular distribution into random noise is comparatively straightforward and can be accomplished by continually adding Gaussian noise, as depicted in Fig.
 2.
-Diffusion model is inspired by this process and realizes the generation of target data by learning the reverse denoising process of this process, which can be traced back to 2015 [42] and became popular after denoising diffusion probabilistic model (DDPM) was published in 2020 [15].
+The concept of diffusion models is inspired by this process and involves learning the reverse denoising process, which dates back to 2015 [42] and gained popularity following the publication of the denoising diffusion probabilistic model (DDPM) in 2020 [15].
 II-A Denoising Diffusion Probabilistic Model (DDPM)
@@
 III-A RS Image Generation
-As one of the most impressive deep generative models, diffusion models are expected to synthesize realistic RS images from existing images or given textual descriptions to support the development of various RS applications.
-According to the data sources, these image generation methods can be mainly divided into two categories: text-to-image generation and image-to-image generation.
+As one of the most impressive deep generative models, diffusion models are expected to synthesize realistic RS images from existing images or given textual descriptions to support the development of various RS applications, as an alternative to autoregressive models [49].
+According to the data sources, these image-generation methods can be mainly divided into two categories: text-to-image generation and image-to-image generation.
 III-A1 Text-to-Image
-Over the past two years, numerous text-to-image diffusion models have come out in computer vision [49, 47, 48, 16], especially the Stable Diffusion (SD) [48] model that has been widely adopted since its release [50, 51, 52].
-However, its success mainly depends on training with billions of text-image pairs from the internet [53], which makes it difficult to extend to the field of RS since such vast and diverse RS datasets are not readily available.
+Over the past two years, numerous text-to-image diffusion models have come out in computer vision [47, 48, 16], especially the Stable Diffusion (SD) [48] model that has been widely adopted since its release [50, 51, 52].
+However, its success mainly depends on training with billions of text-image pairs from the internet [53], which makes it difficult to extend to the field of RS since such vast and diverse RS datasets are not readily available, as mentioned in [49].
 To address this problem, a straightforward idea is to produce trainable RS text-image pairs.
@@
 [104] first extracted features from the cloudy image and noise level, and then inputs the extracted spatial and temporal features into the diffusion model as control conditions.
-Instead of using more control conditions, they trained the model in an supervised manner with cloud-free images to further improve the quality of the reconstructed images.
+Instead of using more controlled conditions, they trained the model in a supervised manner with cloud-free images to further improve the quality of the reconstructed images.
 Moreover, Zhao et al.
@@
 Due to the inherent constraints of imaging technology and environmental conditions, RS images are always accompanied by various noise [106].
-In essence, the learning process of diffusion model is equal to the denoising process [15], which makes it possible to achieve superior denoising performance in the context of RS.
+In essence, the learning process of the diffusion model is equal to the denoising process [15], which makes it possible to achieve superior denoising performance in the context of RS.
 Different modalities of RS images confront different noise challenges.
@@
 However, the RS images always contain diverse and complex scenes, increasing the difficulty of accurate classification.
-Given that diffusion model can learn and simulate complex data distributions better than other deep learning models, researchers are trying to use it for the RS landcover classification.
+Given that the diffusion model can learn and simulate complex data distributions better than other deep learning models, researchers are trying to use it for the RS landcover classification.
 The first application of the diffusion model to this task is presented in [118], based on the conditional diffusion model [44].
@@
 III-C3 Climate Prediction
-Climate prediction is another application for diffusion model in the field of RS, which is a complex systematic project that requires the integration of multiple variables, including cloud amount, cyclone distribution, and water vapor.
+Climate prediction is another application for the diffusion model in the field of RS, which is a complex systematic project that requires the integration of multiple variables, including cloud amount, cyclone distribution, and water vapor.
 Therefore, its solution often consists of multiple diffusion models, enabling fine-grained, stepwise processing of these high-dimensional data.
@@
 Apart from the classic RS interpretation tasks reviewed above, there are some other tasks that may not fall into the above categories [138, 139, 140].
-One of such tasks is height estimation, which aims at providing pixel-wise height information of surface features (e.g., buildings, trees, terrain, etc.) to generate 3D models of surface scenes.
+One such task is height estimation, which aims at providing pixel-wise height information of surface features (e.g., buildings, trees, terrain, etc.) to generate 3D models of surface scenes.
 Diffusion model has been reported to be a promising solution for surface estimation [138].